OpenStudy (anonymous):

Show that y = 1/x is concave up for x > 0 and concave down for x < 0

OpenStudy (anonymous):

Would I start with: y = x^-1 y' = x^-2 y'' x ^-3?

OpenStudy (anonymous):

just sketch it...

OpenStudy (anonymous):

or take the lim as x--->pi 0\[\lim_{x \rightarrow o^{+}} \frac{ 1 }{ x } = + \infty \] \[\lim_{x \rightarrow 0^{-}} \frac{ 1 }{ x } = -\infty\]

OpenStudy (anonymous):

y ' ' = -x^(-2) y ' ' = 2x^(-3) = 2/(x^3) It is clear when x < 0, 2/x^3 is negative and when x > 0, 2/(x^3) is a positive number. Therefore, it is concave up when x > 0, and concave down when x < 0 Keep in mind that if y = x^(-1), then y ' ' = -x^(-2)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok